OpenStudy (aaronandyson):

log 20 + log 50 - 2log 5 - 2log2

OpenStudy (anonymous):

\[\log(a \times b) = \log (a) + \log(b)\]

OpenStudy (aaronandyson):

log 20 times 50 times 4 over 25?

OpenStudy (anonymous):

Break \(20\) with hammer into two parts..

OpenStudy (aaronandyson):

?

OpenStudy (anonymous):

20 = 4 times 5..

OpenStudy (aaronandyson):

oh i thought you said 10 and 10

OpenStudy (anonymous):

\[\log(20) = \log(4 \times 5)\] Now use that formula..

OpenStudy (anonymous):

One more formula you will need here: \[\log(a^n) = n \cdot \log(a)\]

syed98 (syedmohammed98):



syed98 (syedmohammed98):

is it correct @ eta?

OpenStudy (aaronandyson):

It is

OpenStudy (anonymous):

\[\log(4 \times 5) = \log(4 ) + \log(5) = \log(2^2) + \log(5) = 2 \log(2) + \log(5)\]

OpenStudy (anonymous):

Yes, you can go that way too, @SyedMohammed98

OpenStudy (anonymous):

See, in my method, 2 log(2) and -2log(2) will cancel out..

OpenStudy (aaronandyson):

@eta thanks for confusing me.

OpenStudy (anonymous):

@SyedMohammed98 either you proceed or let me do this once, otherwise I know, asker will get confused for sure. I let you do first..@SyedMohammed98, carry on..

syed98 (syedmohammed98):

i will just tell him the formula's log(a) + log(b) = log(a*b) so log(20) + log(50) = log (20*50) log(1000) now the other one which is log(a) - log(b) = log(a/b) so log(5)/)log2) take 2 common so it will be 2 ( (log5)/(log2) ) so if u join both u will get log(1000) - 2log( (log5/log2 )

syed98 (syedmohammed98):

can u continue from here @eta ...or u want to start ur method....

OpenStudy (anonymous):

That will be log(5/2) only.. Yes syed?

OpenStudy (anonymous):

\[2 \log(5) - 2 \log(2) = 2(\log(5) - \log(2) ) = 2 \log(\frac{5}{2})\]

OpenStudy (anonymous):

I think my earlier method was good..

OpenStudy (anonymous):

\[\log(20) = 2 \log(2) + \log(5)\]

OpenStudy (anonymous):

\[\log(50) = \log( 25 \times 2) = \log(25) + \log(2) = \log(5^2) + \log(2) = 2 \log(5) + \log(2)\]

OpenStudy (anonymous):

So, you are left with : \[\implies \log(5) + \log(2) \implies \color{green}{\log(10)}\]

OpenStudy (solomonzelman):

\(\large\color{blue}{ \displaystyle {\rm Log}(20)+ {\rm Log}(50)-2{\rm Log}(5)- 2{\rm Log}(2)}\) \(\large\color{blue}{ \displaystyle {\rm Log}(2)+{\rm Log}(10)+ {\rm Log}(5)+{\rm Log}(10)-2{\rm Log}(5)- 2{\rm Log}(2)}\) \(\large\color{blue}{ \displaystyle 2+{\rm Log}(2)+ {\rm Log}(5)-2{\rm Log}(5)- 2{\rm Log}(2)}\) \(\large\color{blue}{ \displaystyle 2-{\rm Log}(5)- {\rm Log}(2)}\) \(\large\color{blue}{ \displaystyle 2-{\rm Log}(5/2)}\) \(\large\color{blue}{ \displaystyle 2-{\rm Log}(10/4)}\) \(\large\color{blue}{ \displaystyle 2+{\rm Log}(4/10)}\) \(\large\color{blue}{ \displaystyle 2+{\rm Log}(4)-{\rm Log}(10)}\) \(\large\color{blue}{ \displaystyle 1+{\rm Log}(4)}\)

OpenStudy (anonymous):

That is: \[\log_a (a) = 1 \\ \log_{10}(10) = 1\]

OpenStudy (solomonzelman):

eta, for the first line except a=1 or a=0

OpenStudy (anonymous):

But our answer is not matching..

OpenStudy (anonymous):

May be, I have left something..

OpenStudy (solomonzelman):

or maybe I did

OpenStudy (anonymous):

Yes, you must multiply them instead of divide..

OpenStudy (solomonzelman):

\(\large\color{blue}{ \displaystyle 2-{\rm Log}(5)-{\rm Log}(2)}\) \(\large\color{blue}{ \displaystyle 2-\left({\rm Log}(5)+{\rm Log}(2)\right)}\) there is my mistake

OpenStudy (anonymous):

\(2-1 =1\)

OpenStudy (solomonzelman):

\(\large\color{blue}{ \displaystyle 2-\left({\rm Log}(5)+{\rm Log}(2)\right)}\) \(\large\color{blue}{ \displaystyle 2-\left({\rm Log}(10)\right)}\) ypu it is 1

OpenStudy (solomonzelman):

yup*

OpenStudy (aaronandyson):

i am so much confused.

OpenStudy (freckles):

you can also do it this way if it has a pos put it on top if it has a neg put it on bottom this is what I mean \[\log(\frac{20 \cdot 50}{5^2 \cdot 2^2})\]

OpenStudy (freckles):

\[\log(\frac{20}{2^2} \cdot \frac{50}{5^2})\]

OpenStudy (freckles):

you know 20/4=5 and 50/25=2

OpenStudy (freckles):

so you have log(5*2)

OpenStudy (freckles):

can you simplify log(5*2) further

OpenStudy (aaronandyson):

log 10 which is 1

OpenStudy (freckles):

yep which is what @SolomonZelman got above somewhere

OpenStudy (freckles):

all I did was apply the following properties log(a)+log(b)=log(ab) log(m)-log(n)=log(m/n) log(m^r)=r*log(m) and finally you applied log_a(a)=1

OpenStudy (freckles):

\[\text{ Say you have something like this } \\ \log(4)-\log(3)-\log(2)+\log(4)-\log(6) \\ \text{ you can do that whole thingy I mentioned earlier } \\ \log (\frac{4 \cdot 4 }{3 \cdot 2 \cdot 6})\] noticed if the log had a + in front of it the number went on the top inside noticed if the log had a - in front of it the number went on the bottom inside

OpenStudy (aaronandyson):

THANKS!